# Application of the FRADO model of BLR formation to the Seyfert galaxy NGC 5548 and the first step toward determining the Hubble constant

### v2

@@ 4.3 Exemplary shape of the response function
We compute the disk response function and the combined disk-BLR response function using equations 5 and 7, respectively, and present them for nine different wavelengths in the upper and lower panels of Figure 5. As expected, the response function exhibits a bimodal shape, where the first peak corresponds to direct emission from the disk, while the second peak arises from delayed reprocessing by the BLR. This bimodality, driven by the BLR contribution, is clearly evident in the plot. Moreover, the overall time delay associated with the combined response function is significantly longer than that of the disk alone, consistent with the findings of Netzer (2022).
@@ 4.4 Time delays in representative models
This paper serves as a pilot study, focusing on three distinct global setups rather than exploring the full parameter space. In this section, we assume that the source distance can be determined from its redshift, enabling a unique conversion of bolometric luminosities to fluxes. We define these setups as Model A, Model B, and Model C, each characterized by different global parameters. In Models A and C, the black hole mass is derived from observational data, whereas in Model B, it is adopted from the corresponding model reference. A detailed discussion of each model follows in the subsequent sections.
@@ 4.4.1 Model A
In Model A, we minimize the number of free parameters by relying on the basic parameter values from the literature, including the black hole mass, Eddington rate, color correction, and viewing angle (see Table 1). Specifically, the adopted color correction of 2.4 comes from the studies of Kammoun et al. (2021b, 2023). The BLR model is then directly determined based on the assumed black hole mass, Eddington rate, and a fixed metallicity of 5. The resulting predictions are presented in Figure 6. Below, we outline the key advantages and limitations of this model.
@@
The discrepancy between the observed and modeled SED is substantial, differing by a factor of several, and cannot be resolved by adjusting any model parameters. Additionally, in Model A, the time delay is primarily driven by the disk response, with the BLR contribution appearing only near the Balmer edgecontrary to the findings of Netzer (2022). This discrepancy arises because the disk contribution is underestimated due to the fixed color correction being set too high. While this increases the time delay by pushing the fixed-color emission outward, it simultaneously shifts the disk spectrum toward the far-UV. Consequently, we do not include the SED fit from Model A in the minimization, as further adjustments would not improve the fit.
@@ 4.4.2 Model B
In Model B, we adopt the framework of Kubota & Done (2018), where the innermost part of the accretion disk is replaced by a hot corona. In our setup, the standard accretion disk extends from 151 to 282 , and we do not apply any color correction since the disk in this region remains cold. The global system parameters and the inclination angle are consistently taken from Kubota & Done (2018). The free parameters in the model were the BLR contribution, , and the level of starlight. The corona height was fixed at the value of the inner radius of the corona.
The results of this model are presented in Figure 7. As in model A, we employ the same BLR response function (see Section 4.1). However, this approach may introduce minor inconsistencies due to slight differences in the black hole mass and accretion rate. While the time delay fit remains broadly satisfactory, the SED is again not well reproduced. The reduced inner disk radius results in a steep decline in the disk spectrum at shorter wavelengths.
In this configuration, the total time delay is predominantly influenced by the BLR, with only a minimal contribution from the accretion disk. Consequently, the contribution from BLR to the time delay is significantly higher than in model A. Additionally, the amount of starlight in this case is erg s-1 cm-2 Å-1, less than corresponding value in model A. The favored corona height is found to be 43 , consistent with the hot corona size reported in Kubota & Done (2018).
@@ 4.4.3 Model C
In this model, we follow a similar approach to Model B but with considerable modifications. We allow for the variation of the radii dividing the zones. We assume that the cold outer disk is large. We also vary the warm corona parameters as well as the Eddington rate, to achieve the best fit to the broad band spectrum (including X-rays) and to time delays.
The recovered time delays from this model, along with the fitted SEDs, are displayed in the top, middle, and bottom panels of Figure 8, respectively. The Balmer edge is well reproduced in the lag-spectrum. Regarding the Paschen edge, which exhibits a pronounced dip, we do not expect this feature to be fully reproduced. As shown in Pandey et al. (2023) (see their Fig. 4), the ratio of the Balmer to Paschen jump consistently shows that the Balmer jump is usually more prominent.
Additionally, the broad-band SED is now accurately represented, with the starlight continuum at erg s-1 cm-2 Å-1, which is lower than the value of erg s-1 cm-2 Å-1 reported by Mehdipour et al. (2015). f We show only the uv-optical part here for better exposure of the spectral details in the bottom panel. The disk contribution to the time delay in this model falls between the contributions in Model A and Model B. Overall, the time delay is reproduced much more accurately compared to the other two models, providing a better fit to the observed data. The contribution of the corona to the spectrum is dominating the part of the spectrum below Å.
In conclusion, this model offers a significant improvement in both time delay and SED fitting, with a more realistic representation of the BLR contribution to the observed continuum time delays and a better overall alignment with the observed spectral features. However, challenges remain in fully capturing the Paschen edge, indicating areas for further refinement in future modeling efforts.
@@ 5.1 Analytical estimate of
where is the time delay measured at the wavelength , z is the source redshift, is the measured flux at the frequency corresponding to the wavelength , and is the parameter which can take value from 3 to 4. is the viewing angle to the disk.
We can test the model using this formula, but it is important to use the disk delay rather than the total combined disk plus BLR time delay. The disk delay from Model C is 0.880 days at . We convert the flux from Lu et al. (2022), erg s-1cm-2Å-1, to = erg s-1cm-2Hz-1. Assuming an inclination angle of 40 degrees for the source and , we derive of 47.33 km s-1Mpc-1. This represents a considerable improvement over the value of km s-1 Mpc-1 obtained by Cackett et al. (2007) for the same source. The difference is largely due to the fact that we are using only the disk delay. If we had used the total time delay (disk + BLR) of 1.886 days, we would have obtained a much lower value of km s-1 Mpc-1, closer to the result from Cackett et al. (2007).
@@ 5.2 Direct method to measure the distance to the source and from combined spectral and time delay fitting
The Model C in the version described in Section 4.4.3 is based on the known distance which in turn constraints the accretion rate and the bolometric luminosity. Therefore, if we indeed want to measure the distance to the source we cannot rely on quantities that were determined knowing this distance. We thus must allow the Eddington rate to cover a broad range, reflecting the possibly broad range of the luminosity distance of the source. However, we still fix the black hole mass, as it was derived from the spectral fitting (FWHM of the H line) and time delay measurement, so it is not influenced by the knowledge of the luminosity distance.
In this Section, we present a series of computations using a grid of pre-selected luminosity distances. For each luminosity distance, we fit several model parameters by minimizing the value. The parameters we vary when fitting the spectrum include the Eddington ratio (), the cold disk inner radius (), the height of the corona (), the optical depth of the warm corona (), the temperature of the hot corona (), the covering factor of the BLR (), and the normalization of the starlight (). In order to illustrate the relative role of the delay fitting and spectral fitting, we first show results when only one type of data is used (delay or spectra), and only then we show the importance of the combined fitting.
In experiments where only the time delay is fitted, we vary only the first four parameters, keeping the remaining three fixed at the values determined by the spectral best fit, as they do not significantly influence the time delay. In the global fitting method, where both the spectra and the time delays are fitted simultaneously, we also freeze the last three parameters at their best-fit values from the spectral fitting alone. This approach is adopted because spectral fitting is computationally fast, while fitting the time delay is considerably slower, taking over an hour per model. Looking ahead, we aim to accelerate the time delay fitting code or develop a table model to improve computational efficiency.
@@
The dependence of the BLR shape on the incident flux is not monotonic. However, the changes are relatively small, as the range of the considered flux amplitude variation is limited. It is important to emphasize that we do not modify the spectral shape of the incident radiation, as it was derived from observations rather than the model. During spectral fitting, we construct the final model spectrum based on absolute luminosity and compare it to the observed flux, taking into account the adopted distance
@@
To illustrate the roles of time delay fitting and spectral fitting, we first show the results of fitting these two types of observational data separately. The values in the luminosity distance space for the SED-only fit are shown in the upper panel of Figure 10. The fit is not very sensitive to the adopted value of the luminosity distance, . The key parameter that changes with is the accretion rate. The formal minimum of occurs at Mpc, with the Eddington ratio varying from at Mpc to at Mpc. Along with this change in the accretion rate, the corona height decreases from 22 to 12 . Additionally, the covering factor of the BLR, , decreases from 0.42 to 0.30. The stellar contribution also decreases by a factor of 4 as increases. Despite these variations, the optical depth of the hot corona remains unchanged at 21.5, while the electron temperature increases slightly for the two highest values of , staying otherwise at K. Finally, the reduced is smaller than 1, suggesting that the adopted errors can be overestimated.
Complementarily, we also fit only the delay data set in the lag-spectrum, with the corresponding sequence of shown in the lower panel of Figure 10. In this case, the values of the reduced are significantly larger. The dependence of on is relatively weak, with the minimum occurring at Mpc. This sequence exhibits very different properties compared to the spectral sequence. Notably, the Eddington ratio remains constant at 0.015, independent of , since there is no direct coupling between the delay, the distance, and the source luminosity. The weak variation in the irradiation flux does not lead to a substantial change in the model structure.
Therefore, the combined fitting of both the SED and the time lag-spectrum is essential. Spectral fitting enforces a change in the best-fit Eddington ratio with , which, in turn, affects the delay fitting. A small Eddington ratio implies a smaller source size and shorter time delays as expected from standard disk. The results of the combined fitting are shown in the upper panel of Figure 11 in luminosity distance space. Now, in the – plane, each point in the sequence corresponds to the same Eddington ratio and other parameters in both the time delay and spectral fitting. The minimum occurs at Mpc in our grid, though the minimum is relatively shallow. The best fit for both the spectrum and the time delay at this grid is also presented in Figure 8.
The sequence shares some properties with the previously discussed separate solutions. The accretion rate predictably increases along the sequence, while the covering factor decreases. The results are weakly sensitive to the lamp height, which remains close to , as well as to the inner radius of the cold disk, which is located at in all fits.
This can be directly translated into the Hubble constant using the expression , given that the redshift to the source is small. Thus, the same result, but as a function of the implied , is shown in the lower panel of Figure 11. The distance of Mpc corresponds to a Hubble constant of 79.8 km s-1 Mpc-1. Since the minimum is shallow, the error is large, leading to our constraint of .
@@ 6 Discussion
The model itself involves very few parameters. The black hole mass has been fixed, and we have not yet tested its dependence on the results. However, it is worth noting that the black hole mass can be considered independent of the redshift when determined from the line shape. The key parameters – Eddington ratio, BLR contribution, and cold disk inner radius – are fitted using the observed mean spectrum and the continuum time delays measured in the continuum–RM campaign by Fausnaugh et al. (2016). The best fit favors a redshift-independent distance of 62 Mpc to the source. This distance, when converted to the Hubble constant, implies a value of = km s-1 Mpc-1. Additionally, the model successfully reproduces the H time delay. One hidden parameter in the FRADO model is the dust sublimation temperature, , which we adopted as 1500 K. Variations in this temperature would directly influence the favored distance, roughly following the relationship .
The multi-wavelength time delays in NGC 5548 have also been modeled in other studies. The original work by Fausnaugh et al. (2016), which provided the delay measurements, employed only power-law parametric modeling. More recently, Kammoun et al. (2023) fitted the NGC 5548 time delay using a new model that incorporates relativistic effects, such as the X-ray reflection component and albedo, while also exploring the influence of black hole spin and corona height. Their SED fitting yielded a higher Eddington ratio (0.05) than our solution, along with a higher black hole mass, resulting in a significantly greater bolometric luminosity for the source compared to our model. However, we cannot analyze this discrepancy in detail, as their best-fit SED was not explicitly presented. While Kammoun et al. (2023) strongly advocate for the inclusion of color correction, they do not account for the BLR contribution in their model. Notably, in both of these studies, the distance to NGC 5548 was determined using redshift measurements.
@@
Taken together, these comparisons highlight the unique strengths of our approach. Unlike previous studies that either relied on power-law fitting, incorporated relativistic effects without addressing the BLR contribution, or did not pursue an independent distance determination, our method integrates spectral fitting and time delay modeling within a self-consistent framework. By leveraging the FRADO model, we obtained a redshift-independent distance to NGC 5548. While further refinements–such as improved continuum modeling and broader sample studies–are needed to reduce uncertainties, our results demonstrate the potential of this methodology in offering an alternative approach to cosmic distance measurements.

### v3

@@ 4.3 Exemplary shape of the response function
After deriving the BLR response function using FRADO model and the emissivity profile from CLOUDY for NGC 5548, we compute the disk response function and the combined disk-BLR response function using equations 6 and 3.6, respectively. These results are presented for nine different wavelengths in the upper and bottom panels of Figure 5. The disk response function (upper) displays a single-peaked structure, whereas the combined response function (bottom) exhibits a bimodal shape, as expected, with the first peak corresponding to direct emission from the disk, and the second arises from delayed reprocessing by the BLR. Moreover, the overall time delay associated with the combined response function is significantly longer than that of the disk alone, consistent with the findings of Netzer (2022).
@@ 4.4 Time delays in representative models
This paper serves as a pilot study, focusing on three distinct global setups rather than exploring the full parameter space. In this section, we assume that the source distance can be determined from its redshift, enabling a unique conversion of bolometric luminosities to fluxes. We define these setups as Model A, Model B, and Model C, which were shortly introduced in Section 2. Each setup is characterized by different global parameters. In all cases was fixed (in Models A and C, the is derived from observational data (Netzer, 2022)), whereas in Model B, it is adopted from the corresponding model reference (Kubota & Done, 2018). All models are supplemented with the BLR components. A detailed discussion of each model follows in the subsequent sections.
@@ 4.4.1 Model A
To account for disk atmosphere effects, Model A employs a standard accretion disk framework with a fixed color correction. It uses well-established literature values for and Eddington ratio, and does not include the warm corona component. With only a few free parameters, the model incorporates the BLR contribution alongside the accretion disk emission to fit the observed continuum time delays and spectral energy distribution. We minimize the number of free parameters by relying on the basic parameter values from the literature, including the , Eddington rate, color correction, and viewing angle (see Table 1 for the adopted values). Specifically, the adopted color correction of 2.4 comes from the studies of Kammoun et al. (2021b, 2023). The model does not include the warm corona, and the cold disk extends down to the ISCO (6 ). The BLR model is then directly determined based on the assumed , Eddington rate, and a fixed metallicity of 5, but the covering factor is left as a free parameter of the model. The other key parameter is the corona height. The resulting predictions are presented in Figure 6.
The model reproduces the time delay quite well, the Balmer edge at Å is well reproduced, the region of the Paschen jump is not so well fitted. The covering factor and the corona height optimizing the fit of the time delay are given in Table 1.
Below, we outline the key advantages and limitations of this model.
@@
In Model A, the discrepancy between the observed and modeled SED is substantial in the blue part, and cannot be resolved by adjusting any model parameters without making the time delay fit worse. The required lamp height lead to a very blue spectrum in the UV due to the color correction applied. The time delay is primarily driven by the disk response, with the BLR contribution appearing only near the Balmer edgecontrary to the findings of Netzer (2022). This discrepancy arises because the disk contribution is underestimated due to the fixed color correction being set too high. While this increases the time delay by pushing the fixed-color emission outward, it simultaneously shifts the disk spectrum toward the far-UV. Consequently, we do not include the SED fit from Model A in the minimization, as further adjustments would not improve the fit.
@@ 4.4.2 Model B
In Model B, we adopt the framework of Kubota & Done (2018), where the innermost part of the accretion disk is replaced by a hot corona. In our setup, the standard accretion disk extends from 151 to 282 , and we do not apply any color correction since the disk in this region remains cold. The global system parameters and the inclination angle are consistently taken from Kubota & Done (2018). The free parameters in the model are the BLR contribution, , and the level of starlight. The corona height is fixed at the value of the inner radius of the warm corona, i.e., 43 .
The results of this model are presented in Figure 7. As in Model A, we employ the same BLR response function (see Section 4.1). However, this approach may introduce minor inconsistencies due to slight differences in and accretion rate. While the time delay fit remains broadly satisfactory, the SED is again not well reproduced. The reduced inner disk radius results in a steep decline in the disk spectrum at shorter wavelengths.
In this configuration, the total time delay is predominantly influenced by the BLR, with only a minimal contribution from the accretion disk. Consequently, the contribution from BLR to the time delay is significantly higher than in Model A. Additionally, the amount of starlight in this case is erg s-1 cm-2 Å-1, less than corresponding value in Model A.
@@
The modeled time-delay spectrum is predominantly influenced by the BLR, with negligible contribution from the accretion disk, in contrast to observed AGN time-delay spectra that exhibit significant contributions from both the accretion disk and the BLR (Mandal et al., 2025).
@@ 4.4.3 Model C
| Wavelength | ||
| (days) | (days) | |
Note. Columns are (1) pivot-wavelength of the filters in the observed-frame, (2) measured inter-band delays from Fausnaugh et al. (2016) in the rest-frame, (3) recovered inter-band delays from Model C. The inter-band delays are with respect to HST 1367 Å.
This model represents a generalization of Model B. Rather than fixing certain parameters to the values derived by Kubota & Done (2018), we treat them as free parameters and allow the data to constrain them through fitting. First, we allow for the variation of the radii dividing the zones of hot corona, warm corona and outer disk. We assume that the outer radius of the cold disk is large. Then, we vary the warm corona parameters (optical depth, temperature) as well as the Eddington rate, to achieve the best fit to the broad band spectrum (including X-rays), and to time delays. Model parameters are listed in Table 1.
The time delays recovered from this model, along with the observed inter-band delays from Fausnaugh et al. (2016), are presented in Table 2. The corresponding fitted SEDs, are displayed in the upper, middle, and bottom panels of Figure 8, respectively. The Balmer edge is well reproduced in the lag-spectrum. Paschen edge is expected to be less prominent, as shown by Pandey et al. (2023) at the basis of CLOUDY simulations.
Additionally, the broad-band SED is now well represented, with the starlight continuum of erg s-1 cm-2 Å-1, which is nearly identical to the value of erg s-1 cm-2 Å-1 reported by Mehdipour et al. (2015). The middle panel of the figure provides a detailed view of the optical/UV portion of the spectrum, consistent with previous presentations. Furthermore, the bottom panel displays the full X-ray-UV-optical spectral range on a logarithmic scale, allowing for clearer visualization of the broad-band fit. In this representation, both the hard X-ray component and the contribution of the warm corona to the X-ray band are clearly visible. This contrasts with Model A, which lacked a warm corona and therefore failed to reproduce the observed soft X-ray data. In comparison, Model B, by construction from Kubota & Done (2018), is consistent with the X-ray band and thus its X-ray spectral range is not plotted in this figure. However, it is important to note that Kubota & Done (2018) did not fit the spectroscopic optical data, and as a result, the discrepancy between their model and the actual optical observations was not apparent in their analysis.
The disk contribution to the time delay in this model falls between the contributions in Model A and Model B. Overall, the time delay is reproduced much more accurately compared to the other two models, providing a better fit to the observed data. The contribution of the corona to the spectrum is dominating the part of the spectrum below Å.
@@
In conclusion, this model offers a significant improvement in both time delay and SED fitting, with a more realistic representation of the BLR contribution to the observed continuum time delays and a better overall alignment with the observed spectral features.
@@ 5.1 Analytical estimate of
where is the time delay measured at the wavelength , is the source redshift, is the measured flux at the frequency corresponding to the wavelength , and is the parameter which is used to account for systematic discrepancies due to conversion of annulus temperature to the corresponding wavelength at a given radius. According to Wien’s law, (Netzer, 2022); however, under the assumption of a flux-weighted radius, is typically taken to be 2.49 (Edelson et al., 2024). In our analysis, considering the AGN accretion disk SED as a multicolored blackbody spectrum, we adopt .
We can test the model using this formula, but it is important to use the disk delay rather than the total combined disk plus BLR time delay. The disk delay from Model C is days at Å relative to the X-ray source. For a corona height of , this delay becomes days (i.e., days), and this value is therefore used in equation (11). We convert the flux from Lu et al. (2022), erg s-1cm-2Å-1, to = erg s-1cm-2Hz-1. Assuming an inclination angle of degrees for the source, we derive of 47.54 km s-1Mpc-1. This represents a considerable improvement over the value of km s-1 Mpc-1 obtained by Cackett et al. (2007) for the same source. The difference is largely due to the fact that we are using only the disk delay. If we had used the total time delay (disk + BLR) of 2.00 days, we would have obtained a much lower value of km s-1 Mpc-1, closer to the result from Cackett et al. (2007).
However, adopting yields a derived of 25.27 km s-1 Mpc-1 for the disk delay, or 11.34 km s-1 Mpc-1 when including both the disk+BLR delay. Since there is no universally accepted value of , equation 11 inherently carries additional uncertainty associated with this parameter.
@@ 5.2 Direct method to measure the distance to the source and from combined spectral and time delay fitting
The Model C in the version described in Section 4.4.3 is based on the known distance which in turn constraints the accretion rate and the bolometric luminosity of the source. Therefore, if we indeed want to measure the distance to the source we cannot rely on quantities that were determined knowing its distance. We thus must allow the Eddington rate to cover a broad range, reflecting the possibly broad range of the luminosity distance of the source. However, we keep fixed, as it is derived from spectral fittingusing the FWHM/line dispersion () of the H line and the time delay measured from RM, making it independent of the luminosity distance. We also fix the viewing angle.
In this Section, we present a series of computations using a grid of pre-selected luminosity distances. For each luminosity distance, we fit several model parameters by minimizing the value. The parameters we vary when fitting the spectrum include the Eddington ratio (), the cold disk inner radius (), the height of the corona (), the optical depth of the warm corona (), the temperature of the hot corona (), the covering factor of the BLR (), and the normalization of the starlight ().
@@
The dependence of the BLR shape on the incident flux is not monotonic. However, the changes are relatively small, as the range of the considered flux amplitude variation is limited. It is important to emphasize that we do not modify the spectral shape of the incident radiation, as it was derived from observations rather than the model. During spectral fitting, we construct the final model spectrum based on absolute luminosity and compare it to the observed flux, taking into account the adopted distance as
@@
The results of the combined delay and spectra fitting are shown in the upper panel of Figure 10 in luminosity distance space. Now, in the – plane, each point in the sequence corresponds to the same Eddington ratio and other parameters in both the time delay and spectral fitting. The minimum occurs at Mpc in our grid, though the minimum is relatively shallow. The best fits for both the spectrum and the time delay at this grid are also presented in Figure 8.
The accretion rate predictably increases along the sequence, from 0.008 to 0.038, while the covering factor remain roughly the same, at %. The results are weakly sensitive to the lamp height, which remains close to , as well as to the inner radius of the cold disk, which is located at in all fits.
The value of the luminosity distance can be directly translated into the Hubble constant using the expression , given that the redshift to the source is small. Thus, the same result, but as a function of the implied , is shown in the lower panel of Figure 10. The distance of Mpc corresponds to a Hubble constant of 66.8 km s-1 Mpc-1. Interpolating between the grid points we find the minimum location at 73.96 Mpc. Since the minimum is shallow, the error is large, leading to our constraint of as Mpc. This translates to Hubble constant constraints: .
The error estimate presented here is necessarily approximate, as it is based on a qualitative assessment of the effective number of independent degrees of freedom (dof) in the model. Our dataset includes 17 measurements of time delays and we sample the spectrum in the same number of wavelengths. The model involves 7 free parameters, in addition to 4 parameters that are held fixed: , viewing angle, local cloud density, and column density. Each model realization is computed for a fixed luminosity distance. This setup yields 23 dof. We simply determine the 1 sigma error of the reduced as . We did not perform tests whether all these measurements are statistically independent.
A more robust approach to error estimation would be to use a bootstrap resampling method, which could account for eventual correlations. Unfortunately, our current implementation is not optimized for such computationally intensive analysis, which would require at least 100 realizations to obtain reliable error estimates. Code optimization is underway, as we view this methodology as promising for future studies.
@@
To illustrate the necessity of combining delay and spectral fitting, we also present the results obtained when only one type of data, either delays or spectra is used independently.
In experiments where only the time delay is fitted, we vary four parameters (, , , ), while the other three (, , ) are fixed at the values determined by the spectral best-fit, as these parameters have a negligible impact on the time delay. This strategy is motivated by the fact that spectral fitting is computationally efficient, whereas time delay fitting is considerably more time-consuming. Moreover, when fitting the SED, all parameters must be optimized simultaneously.
The values in the luminosity distance space for the SED-only fit are shown in the upper panel of Figure 11. The key parameter that changes with is the accretion rate. The formal minimum of occurs at Mpc, with the Eddington ratio varying from at Mpc to at Mpc. Along with this change in the accretion rate, the corona height increases from 16.5 to 53.0 . The covering factor of the BLR, , remains roughly the same, . The stellar contribution also decreases by a factor of 4 as increases. Despite these variations, the optical depth of the hot corona remains roughly unchanged at 20, while the electron temperature increases slightly with , from K to K. The reduced is of order of 1, suggesting that the adopted spectral errors of 7% are realistic.
The values for the case when we fit only the delay data set in the lag-spectrum, are shown in the lower panel of Figure 11. The dependence of on is relatively weak, exhibiting an almost monotonic decrease in values with increasing luminosity distance. This sequence exhibits very different properties compared to the spectral sequence. The Eddington ratio remains constant at 0.03, independent of , since there is no direct coupling between the delay, the distance, and the source luminosity. The lamp height is monotonically increasing with . Additionally, the weak variation in the irradiation flux does not lead to a substantial change in the model structure.
Therefore, the combined fitting of both the SED and the time lag-spectrum is essential. Spectral fitting enforces a change in the best-fit Eddington ratio with , which, in turn, affects the delay fitting. A small Eddington ratio implies a smaller source size and shorter time delays as expected from standard disk.
@@ 6 Discussion
The model itself involves very few parameters. has been fixed, and we have not yet tested its dependence on the results. However, it is worth noting that can be considered independent from cosmology since it is based on line kinematic width and the measured time delay. The key parameters – Eddington ratio, BLR contribution, and cold disk inner radius are fitted using the observed mean spectrum and the continuum time delays measured in the continuum–RM campaign by Fausnaugh et al. (2016). The best fit favors a redshift-independent distance of Mpc to the source. This distance, when converted to the Hubble constant, implies a value of = km s-1 Mpc-1. Additionally, the model successfully reproduces the H time delay. One hidden parameter in the FRADO model is the dust sublimation temperature, , which we adopt as 1500 K. Variations in this temperature would directly influence the favored distance, roughly following the relationship .
The multi-wavelength time delays in NGC 5548 have also been modeled in other studies. The original work by Fausnaugh et al. (2016), which provided the delay measurements, employed only power-law parametric modeling. More recently, Kammoun et al. (2023) fitted the NGC 5548 time delay using a new model that incorporates relativistic effects, such as the X-ray reflection component and albedo, while also exploring the influence of black hole spin and corona height. Their SED fitting yielded a higher Eddington ratio (0.05) than our solution, along with a higher , resulting in a significantly greater bolometric luminosity for the source compared to our model. However, we cannot analyze this discrepancy in detail, as their best-fit SED was not explicitly presented. While Kammoun et al. (2023) strongly advocate for the inclusion of color correction, they do not account for the BLR contribution in their model. Notably, in both of these studies, the distance to NGC 5548 was determined using redshift measurements.
@@
Constructing a complete model of the BLR remains a significant challenge, with progress being made along several fronts. For instance, Long & Dexter (2025) analyzed the STORM data for NGC 5548 and, by combining a few representative geometries with table models that span a range of cloud ionization properties, concluded that none of the simplest models adequately reproduce the observed data. This highlights the limitations of overly simplified assumptions. Additionally, parametric approaches, such as those implemented in the BELMAG code (Rosborough et al., 2024, 2025), exemplify the complexity and richness required to capture the full diversity of BLR structures.
Our approach does not yet incorporate a detailed treatment of the ionization levels of individual clouds, as our objective is more modest and thus we do not aim to reproduce all features of the available data for the NGC 5548 source. Instead, we begin by establishing the dynamics based on the FRADO model, which alone enables us to roughly reproduce key observables such as the H time delay, the transfer function, and the mean spectrum (apart from intensities of all lines), as well as the overall time-delay spectrum.
The geometry adopted in our model naturally leads to double-peaked emission lines. This is consistent with the observed double-peaked structure of the H line, particularly in the rms profile (e.g. Lu et al., 2016; Shapovalova et al., 2009; Long & Dexter, 2025). Moreover, other emission lines also exhibit double-peaked profiles (Horne et al., 2021), and the velocity-resolved time delays further support this structure (Long & Dexter, 2025). Our model recovered transfer function similarly displays a double-peaked profile, consistent with an axially symmetric geometry dominated by Keplerian motion. In contrast, Horne et al. (2021) obtained an almost single-peaked transfer function for H, which may indicate the presence of cloud emission anisotropy, an effect already considered by Ferland et al. (1979), as well as possible selective shielding of parts of the BLR by outflowing material located within the region.
There remains considerable room for advancement in the coming years, particularly in refining the physical details of BLR modeling. Furthermore, the potential application of continuum time-delay measurements to cosmology provides an additional strong motivation for continued progress in this area.
Taken together, these comparisons highlight the unique strengths of our approach. Unlike previous studies that either relied on power-law fitting, incorporated relativistic effects without addressing the BLR contribution, or did not pursue an independent distance determination, our method integrates spectral fitting and time delay modeling within a self-consistent framework. By leveraging the FRADO model, we obtain a redshift-independent distance to NGC 5548. While further refinements–such as improved continuum modeling and broader sample studies–are needed to reduce uncertainties, our results demonstrate the potential of this methodology in offering an alternative approach to cosmic distance measurements.